OpenStudy (anonymous):

Please explain how 40Newtons *cos30 = 34.6 newtons

OpenStudy (anonymous):

\[\cos(30^\circ)=\frac{\sqrt{3}}{2}= .866...\] im going to assume 40 *.866 equals about 34.6

OpenStudy (anonymous):

does this make sense?

OpenStudy (anonymous):

yes thanks alot